# AnyV2V: A Tuning-Free Framework For Any Video-to-Video Editing Tasks

### v3

@@ D.4 Safeguards
D.5 Ethical Concerns for Human Evaluation
@@ D.6 New Assets
Appendix E More AnyV2V Showcases
Figure 11: AnyV2V becomes an instruction-based video editing tool when plugged with instruction-guided image editing models like InstructPix2Pix [4]. Prompt used “Turn the couple into robots”, “Turn horse into zebra”, and “Turn the sand into snow”.
Figure 12: The more recent model InstantStyle [48] can seamlessly plug in AnyV2V to perform reference-based style transfer video editing. The reference art image in the bottom left corner is used to retrieve the first image edit.
Figure 13: AnyV2V can perform subject-driven video editing with a single image reference, using a subject-driven image editing model like AnyDoor [9]. The subject image in the bottom left corner is used to retrieve the first image edit.

### v4

@@ C.4 Safeguards
C.5 Ethical Concerns for Human Evaluation
@@ C.6 New Assets
Appendix D More AnyV2V Showcases
Figure 11: AnyV2V becomes an instruction-based video editing tool when plugged with instruction-guided image editing models like InstructPix2Pix (Brooks et al., 2023). Prompt used “Turn the couple into robots”, “Turn horse into zebra”, and “Turn the sand into snow”.
Figure 12: The more recent model InstantStyle (Wang et al., 2024a) can seamlessly plug in AnyV2V to perform reference-based style transfer video editing. The reference art image in the bottom left corner is used to retrieve the first image edit.
Figure 13: AnyV2V can perform subject-driven video editing with a single image reference, using a subject-driven image editing model like AnyDoor (Chen et al., 2023c). The subject image in the bottom left corner is used to retrieve the first image edit.